### v2

# Understanding Chain-of-Thought in LLMs through Information Theory

Abstract
Large Language Models (LLMs) have shown impressive performance in complex reasoning tasks through the use of Chain-of-Thought (CoT) reasoning, allowing models to break down problems into manageable sub-tasks. However, existing CoT evaluation techniques either require annotated CoT data or fall short in accurately assessing intermediate reasoning steps, leading to high rates of false positives. In this paper, we formalize CoT reasoning in LLMs through an information-theoretic lens. Specifically, our framework quantifies the ‘information-gain’ at each reasoning step, enabling the identification of failure modes in LLMs without the need for expensive annotated datasets. We demonstrate the efficacy of our approach through extensive experiments on toy arithmetic, GSM8K and PRM800k datasets, where it significantly outperforms existing outcome-based methods by providing more accurate insights into model performance on individual subtasks.
[…]
1 Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities across a wide range of tasks, from complex reasoning to code generation [4, 23, 3, 2]. Many of these advances can be attributed to Chain-of-Thought (CoT) reasoning [32, 22, 17], which involves breaking down complex problems into a series of intermediate steps, mirroring human-like reasoning processes. The success of CoT reasoning, particularly in domains such as mathematics, logic, and multi-step decision-making, has led researchers to incorporate CoT-like features directly into model training, i.e. the FLAN family of models [5, 31].
This paper introduces a new formal framework for analyzing CoT in LLMs. We provide a rigorous method grounded in information theory, to evaluate the quality of each step in a model’s reasoning process, thus offering insights beyond simple accuracy metrics to identify areas for improvement.
Previous work in this area has proposed “Process Supervision[18], which requires expensive, human-annotated step-by-step data. While effective, this approach is often impractical due to the high cost and effort of creating large-scale annotated datasets. In turn, alternative methods have recently been proposed, such as outcome reward modelling [11] or the Math-Shepherd [29]. Both these approaches avoid reliance on costly annotated step-wise CoT data by instead modelling the correctness of each step based on the correctness of final outputs. However, as we show later, these methods can be unsound for detecting incorrect reasoning steps and can thus lead to a high false-positive rate in certain scenarios.
To address these shortcomings, we employ an information-theoretic approach, grounded in the following key insight: Each correct step in a reasoning process should provide valuable and relevant information that aids in predicting the final correct outcome. Building on this insight, we develop a framework to quantify the “information-gain” after each sub-task in the reasoning process, without the need for step-by-step annotations. This enables us to detect sub-tasks that fail to contribute meaningful information toward the correct solution, signalling potential errors or irrelevant steps in the model’s reasoning. In addition, we also introduce a practical algorithm to assess LLM performance across various sub-tasks within a Chain-of-Thought (CoT) reasoning process. The key contributions of this paper are as follows:
  1. 1.
    We develop a framework for sequential applications of sub-tasks, e.g. Chain-of-Thought and provide a rigorous language to identify failure modes in LLMs.
  2. 2.
    Based on this framework, we propose a practical algorithm to assess the task-wise performance of models. This yields more granular information about a model’s CoT performance without requiring annotated data for intermediate reasoning steps.
  3. 3.
    We validate our methods on extensive toy data, the GSM8K [6] as well as the PRM800K [18] dataset. Our method effectively identifies failure modes in CoT reasoning, unlike baselines like outcome reward modelling [11] and Math-Shepherd [29], which rely on final accuracy and tend to increase false positives in error detection.
2 Proposed Framework: Setup and Notation
Before diving into our framework, we first provide a high-level overview and notation on how LLM generation will be treated throughout this paper. This will allow us to set the foundation for describing our information-theoretic framework. In particular, following the approach in [10], we view LLMs as abstract execution machines with a natural language interface. From this perspective, prompts are designed to solve specific problems (e.g., mathematical or logical problems), and the LLM processes the information in the prompt to generate an output.
We define a typical prompt as a combination of two parts:
  1. 1.
    An initial state, represented by the random variable $X_{0}\in\mathcal{X}$, encapsulates the prompt-provided information that the LLM processes to derive the queried result.
  2. 2.
    A task $\lambda\in\Upsilon$ (e.g., addition then multiplication) defines how the LLM processes $X_{0}$.
Given the prompt, defined as a tuple $(X_{0},\lambda)$, the state $X_{1}$ represents the result of applying task $\lambda$ to the initial state $X_{0}$. Formally, we denote this using the update mapping $\Lambda:\mathcal{X}\times\Upsilon\rightarrow\mathcal{X}$ which outputs the updated state $X_{1}$ by applying the task $\lambda$ on $X_{0}$, i.e. $X_{1}=\Lambda(X_{0},\lambda)$. This updated state is then used to obtain the final output, denoted by $Y\in\mathcal{X}$, by extracting only the information in $X_{1}$ which is relevant to the queried final answer. This notation defines a prompt that instructs a model to process information drawn from some initial distribution $p(X_{0})$ (e.g., math problems).
We use a simple example to illustrate this notation:
Prompt: “Solve for $z=2\,(u+v)$, $u=12$, $v=13$ (1)
Here, the initial state $x_{0}$ denotes the information $u$ = 12, $v$ = 13”, and $\lambda$ denotes the task of finding $z$ (i.e. addition followed by multiplication). Next, $x_{1}=\Lambda(x_{0},\lambda)$ represents the updated information after correctly performing the addition operation, i.e. $x_{1}$ is $u$ = 12, $v$ = 13 and $z$ = 50”. The final output, $y$, is then obtained by simply extracting the value of $z$ from $x_{1}$, i.e. $z$ = 50”.
Remark Our setup also encapsulates cases with ambiguous (or multiple correct) responses for a given task $\lambda$. In this case, $\Lambda(x_{0},\lambda)$ is a random variable with distribution $p(X_{1}\mid X_{0}=x_{0})$. Therefore, for generality, we treat $\Lambda(x_{0},\lambda)$ as a random variable from now on.
2.1 Compositionality
Many mathematical or logical problems, such as the one in (1), require sequential application of operations. Our notation is also amenable to such problems as it accommodates the composition of tasks.
For example, one way to address prompt (1) involves first adding $u$ and $v$, and next, multiplying the result by $2$ to find $z$. Using our notation, this can be expressed as $\Lambda(x_{0},\lambda_{1}\circ\lambda_{2})$, where $\lambda_{1},\lambda_{2}$ denote the addition and multiplication tasks respectively. The following property allows us to define the application of compositional task $\lambda_{1}\circ\lambda_{2}$:
Definition 2.1.
[…]
Returning to the prompt in (1), Figure 1 shows that the model first computes $u+v$, then multiplies the result by $2$.
Figure 1: Prompt (1) requires compositional application of tasks.
Here, we refer to $X_{1},X_{2}$ as intermediate states and $Y$ is the correct final output. In general, if a problem statement requires sequential application of $T$ sub-tasks, $\lambda=\lambda_{1}\circ\ldots\circ\lambda_{T}$, then the Chain-of-Thought (CoT) reasoning is divided up into $T$ steps, where the $t$’th step is recursively defined as $X_{t}=\Lambda(X_{t-1},\lambda_{t})$ for $t\in\{1,\dots,T\}$. Finally, the overall true output $Y$ is obtained by extracting the queried information from the final state $X_{T}$.
Having established a formal language for the sequential application of tasks, we now turn towards how a task may be divided into such a sequence of intermediate sub-tasks.
2.2 Primitive Tasks
In this subsection, we introduce the notion of primitive tasks which form the basic building blocks of any task. Intuitively, our formulation is reminiscent of ideas from linear algebra, where basis vectors form the basic building blocks of a vector space. In our case, any task $\lambda\in\Upsilon$ can be expressed as a sequence of primitive tasks. This decomposition will allow us to establish which tasks the model could have learned from the training data. For example, if a specific primitive task is not available in the LLM training data, it would be impossible for the model to execute any instructions which involve this primitive task correctly. With this in mind, we now introduce this concept formally:
Definition 2.2 (Primitive tasks).
We say that a set of tasks $\Gamma\subseteq\Upsilon$ is primitive if, for any task $\lambda\in\Upsilon$, there exists a unique subset $\{\lambda_{i}\}_{i=1}^{k}\subseteq\Gamma$ such that $\lambda=\lambda_{1}\circ\dots\circ\lambda_{k}$.
Note that the decomposition is not unique but the set of components is. In some cases, there may exist distinct permutations of primitive tasks which compose to yield the same task as is common in many associative operations. As an example, in the context of mathematical problem-solving, the basic arithmetic operations could be considered primitive. The composition of these primitive tasks allows us to construct extremely complex operations. Just like in linear algebra, we define the span of these tasks as the set obtained by their sequential applications.
Definition 2.3 (Span of tasks).
Let $\Phi\subseteq\Upsilon$ be a set of tasks:
$\textup{Span}(\Phi)=\{\lambda_{1}\circ\ldots\circ\lambda_{k}:\lambda_{i}\in\Phi\textup{ for }1\leq i\leq k,k\in\mathbb{Z}_{>0}\}.$
The set $\textup{Span}(\Phi)$ comprises all the tasks that can be applied by composing sub-tasks in the set $\Phi$. This means that any compositionally consistent update rule $\Lambda$ which is well-defined on the set of tasks $\Phi$ will also be well-defined on $\textup{Span}(\Phi)$. However, this $\Lambda$ may still be ill-defined for any task not in this span. This limitation, known as unidentifiability, defines the boundaries of a model’s inferences.
2.3 Unidentifiability
The unidentifiability of tasks forms a key part of our framework. It directly addresses the fundamental challenge that models, such as LLMs, face when dealing with unseen tasks. If a task $\lambda$ lies outside of $\textup{Span}(\Phi)$, the span of tasks the model has been trained on, then the model cannot be expected to infer or apply it correctly. In other words, the model’s capacity is constrained by the identifiability of tasks within the training set. This notion and formalization of unidentifiability allows us to highlight a critical limitation in the generalization of models: tasks not encountered during training cannot be reliably executed, as they remain beyond the model’s learned task-span. More formally:
Definition 2.4 (Unidentifiability).
A task $\lambda$ is unidentifiable in a set $\Phi\subseteq\Upsilon$ if and only if $\lambda\not\in\textup{Span}(\Phi)$.
Remark   In practice, unidentifiability may depend on the initial state $X_{0}$, i.e. an LLM might accurately perform addition for 2-digit numbers but fail with 10-digit numbers [24]. For more details, see Appendix 8.1.
Building on this framework, we propose an algorithm that integrates unidentifiability with information-theoretic methods to detect CoT reasoning failures.
3 Operationalising Our Framework
This section aims to operationalise the above framework to make inferences regarding the unidentifiability of intermediate sub-tasks in a model’s CoT reasoning process. This would subsequently allow us to detect any sub-task at which a model’s CoT reasoning process starts to diverge from the ground truth, thereby providing insights into how the model can be improved. For example, suppose we are in a setting where the “addition” operation is unidentifiable, then we could further improve the model’s mathematical reasoning by fine-tuning it on the addition operation.
3.1 An information-theoretic perspective
To apply unidentifiability in CoT generations, we introduce a fundamental assumption: each correct CoT step should provide meaningful information aiding the prediction of $Y$. If a step ceases to increase information about $Y$, it indicates an incorrect execution. We formalize this assumption using our notation from the previous section:
Assumption 3.1 (Bayesian network).
Let $\lambda\neq\lambda^{\prime}$ be two operations with primitive decompositions:
$\lambda$ $=\lambda_{1}\circ\dots\lambda_{k-1}\circ\lambda_{k}\circ\dots\circ\lambda_{T}\quad\text{ and }$
$\lambda^{\prime}$ $=\lambda_{1}\circ\dots\lambda_{k-1}\circ\lambda^{\prime}_{k}\circ\dots\circ\lambda^{\prime}_{T^{\prime}},$
where $\lambda^{\prime}_{k}$ is unidentifiable in $\{\lambda_{1},\dots,\lambda_{T}\}$. Then, the intermediate states corresponding to the tasks $\lambda,\lambda^{\prime}$ have the Bayesian network in Figure 2.
[…]
Figure 2: Bayesian network corresponding to Assumption 3.1.
Intuition
The Bayesian network in Figure 2 implies that if we encounter an unidentifiable task ($\lambda^{\prime}_{k}$) at step $k$ of the reasoning path, the future states $X_{i}$ and $X^{\prime}_{j}$ for any $i,j\geq k$ satisfy the conditional independence $X_{i}\perp\!\!\!\perp X^{\prime}_{j}\mid X_{k-1}$. Consequently, once we apply $\lambda^{\prime}_{k}$, the subsequent states along the new reasoning path (in red) add no information regarding the subsequent states or the output of the original path (in green). Hence the figure represents the fact that, for any given input, the output of $\lambda_{k}$ (top fork) contains no information regarding the output of any other primitive task $\lambda^{\prime}_{k}$ (bottom fork).
With our key assumption on the ground-truth CoT process formalized, we now consider the model’s CoT behaviour.
3.2 Task execution in LLMs
To operationalise our framework, we formally distinguish between the model i.e. LLM’s task execution and the ground truth process which arises from following the instructions correctly. To this end, we explicitly define how an LLM interprets a specified task $\lambda$ using $\Lambda^{M}(X_{0},\lambda)$, which is in general distinct from the ground truth update rule $\Lambda(X_{0},\lambda)$.
Here, one option would be to consider the idealised setting where the model learns to perfectly follow some of the primitive tasks available in the training data. However, this may be considered too restrictive since in reality most LLMs do not always follow a “learned” task perfectly. Instead, we consider a much weaker assumption that the model cannot correctly execute a task which is unidentifiable in the training data. Concretely, suppose $\Gamma^{M}\subseteq\Gamma$ denotes the primitive tasks available in the LLM training data. Then, we make the following assumption on LLM’s task execution.
Assumption 3.2 (Task execution in LLMs).
[…]
Intuition
Assumption 3.2 means that for any task which we would like the LLM to apply, the LLM ends up executing some task in $\textup{Span}(\Gamma^{M})$ which the model has been trained on. In other words, the model’s execution is restricted only to the tasks which could be inferred from the training data (i.e. in $\textup{Span}(\Gamma^{M})$). Moreover, this assumption also allows us to encapsulate cases where the model does not follow correct instructions or does not decompose a task correctly.
Before proceeding further with our main result which will allow us to test for the unidentifiability of sub-tasks, we define some notation which we will use from now onwards. Let $\lambda=\lambda_{1}\circ\ldots\circ\lambda_{T}$ denote a primitive decomposition of a task $\lambda$. Then, starting from an initial state $X_{0}$, we denote the model’s intermediate states recursively as:
$X^{M}_{t}\coloneqq\Lambda^{M}(X^{M}_{t-1},\lambda_{t})\quad\textup{and}\quad X^{M}_{0}=X_{0}.$
Moreover, we use $Y^{M}$ to denote the model’s final output. Next, using this notation, we present the conditional independence which must hold if the model encounters an unidentifiable intermediate task along its reasoning path.
Theorem 3.3.
Let $\Gamma^{M}\subseteq\Gamma$ denote the primitive tasks available in the training data. Let $\lambda$ be a task with decomposition $\lambda=\lambda_{1}\circ\ldots\circ\lambda_{T}$. If $\lambda_{k}$ is the first task in the decomposition of $\lambda$ which is unidentifiable in $\Gamma^{M}$ (i.e. $k=\arg\min_{t}\{\lambda_{t}\not\in\textup{Span}(\Gamma^{M})\}$). Then, under Assumptions 3.1 and 3.2, we have that
$Y\perp\!\!\!\perp X^{M}_{j}\mid X^{M}_{j-1}\quad\textup{for all $j\geq k$}.$ (2)
Theorem 3.3 shows that under Assumptions 3.1 and 3.2, when the model encounters an unidentifiable task (i.e. $\lambda_{k}$ in Theorem 3.3) in its Chain-of-Thought reasoning, the model output satisfies the conditional independence in Equation (2). In practice, this means that if at step $k$, a model encounters a reasoning step which is necessary for obtaining the correct answer and is unidentifiable in the training data, the CoT reasoning diverges from the ground truth at this step and every subsequent step adds no additional information regarding the correct final output Y. This ‘information’ can be measured by checking if the model’s confidence about the final output $Y$ increases after each step. This is formalised in the next section.
3.3 Testing for unidentifiability using information-gain
With our framework established, we now describe how to detect unidentifiable sub-tasks using information theory. Following common practice [29, 11], we assume access to a dataset of prompts and correct final answers, derived by applying task $\lambda$, denoted as $\mathcal{D}_{\lambda}\coloneqq\{(x_{0}^{i},y^{i})\}_{i=1}^{n}$. Recall that $X^{M}_{j}$ and $X^{M}_{j-1}$ represent the model’s chain of thought (CoT) reasoning at steps $j$ and $j-1$, respectively. Consequently, each element in the conditional independence statement in Equation (2) can be derived from the data and/or the model.
To this end, we consider the mutual information between $Y$ and $X^{M}_{j}$ conditional on $X^{M}_{j-1}$, denoted by $\mathcal{I}(Y;X^{M}_{j}\mid X^{M}_{j-1})$. This conditional mutual information term intuitively represents the additional information contributed by the $j$’th step of CoT, which is relevant for predicting the ground truth final output $Y$. Therefore, we refer to $\mathcal{I}(Y;X^{M}_{j}\mid X^{M}_{j-1})$ as the information-gain at step $j$.
It follows from Theorem 3.3 that if an LLM encounters a sub-task at step $i$ which is unidentifiable in its training data, no subsequent step should contribute any additional information relevant for predicting $Y$ (i.e. the information-gain should remain 0 after step $i$). If, on the other hand, we observe that $\mathcal{I}\left(Y;X^{M}_{j}\mid X^{M}_{j-1}\right)>0$ for some $j\geq i$, then under Assumptions 3.1 and 3.2, the task $\lambda_{i}$ is not unidentifiable. To estimate the information-gain in practice, we use the following result:
Proposition 3.4.
Let $\mathcal{I}(X;Y\mid Z)$ denote the mutual information between $X$ and $Y$ conditional on $Z$. Then,
$\mathbb{E}[\log p(Y\mid X^{M}_{j})]-\mathbb{E}[\log p(Y\mid X^{M}_{j-1})]=\mathcal{I}\left(Y;X^{M}_{j}\mid X^{M}_{j-1}\right)\geq 0.$ (3)
To estimate the information-gain in (3) using Proposition 3.4, we train a separate LLM, which we refer to as the supervisor model $g_{\textup{sup}}$. This model takes as input the model’s CoT reasoning up to any given intermediate step $t$, $X^{M}_{t}$, and is fine-tuned to directly predict the ground truth final output $Y$. In this way $g_{\textup{sup}}(X^{M}_{t})$ approximates the conditional distribution $p(Y\mid X^{M}_{t})$. Then, the quantity $\mathbb{E}[\log p(Y\mid X^{M}_{j})]$ can be estimated using the negative cross-entropy loss for predicting $Y$, i.e., $\mathbb{E}[\log p(Y\mid X^{M}_{j})]$ is approximately
$\mathbb{E}[\log\hat{p}(Y\mid X^{M}_{j})]=-\mathbb{E}[l_{\textup{CE}}\left(Y,g_{\textup{sup}}(X^{M}_{j})\right)],$
where $l_{\textup{CE}}$ denotes the cross-entropy loss. From this, it follows that
$\underbrace{\mathbb{E}[\log p(Y\mid X^{M}_{j})]-\mathbb{E}[\log p(Y\mid X^{M}_{j-1})]}_{\text{Information-gain}}\approx\mathbb{E}[l_{\textup{CE}}(Y,g_{\textup{sup}}(X^{M}_{j-1}))]-\mathbb{E}[l_{\textup{CE}}(Y,g_{\textup{sup}}(X^{M}_{j}))].$ (4)
Summary    The information-gain (IG) between steps $j$ and $j-1$ reflects how much relevant information step $j$ contributes towards predicting $Y$. If task $\lambda_{j}$ is executed correctly, this gain is positive, as indicated by a decrease in the cross-entropy loss. Conversely, if step $j$ does not provide additional information, the loss remains unchanged. This can be interpreted as the conditional mutual information between $X^{M}_{j}$ and $Y$, conditioned on $X^{M}_{j-1}$. Positive information-gain suggests step $j$ adds new insight about $Y$, while no gain indicates no added information. Training details for the supervisor model are in Appendix 9.1.3.
Remark on sample-wise information-gain     While conditional mutual information provides an aggregate measure of information-gain for a sub-task in a dataset, it may also be desirable to obtain an analogous measure of sub-task correctness for individual CoT instances. This could be useful, for example, in detecting which step is wrong for a given prompt. Our notion of information-gain can be extended to this sample-wise setting, similar to [8], by instead considering the following difference
$\log p(Y\mid X^{M}_{j})-\log p(Y\mid X^{M}_{j-1})\approx l_{\textup{CE}}(Y,g_{\textup{sup}}(X^{M}_{j-1}))-l_{\textup{CE}}(Y,g_{\textup{sup}}(X^{M}_{j})).$ (5)
Intuitively, if step $j$ in the model’s CoT is correct, the model should become more confident in the ground truth output $Y$ being the correct final answer. Therefore, the difference above should be positive. Conversely, if step $j$ is wrong, the model’s confidence regarding $Y$ should not increase, and this difference should be $\leq 0$. From now on, we refer to the difference in (5) as sample-wise information-gain at step $j$.
Remark on O1/R1 style reasoning     Although we present our framework using linear chains-of-thought for clarity, the information-gain metric naturally accommodates the more complex reasoning patterns found in O1/R1-style models. These reasoning models often explore multiple solution paths, backtrack when encountering errors, and dynamically self-correct their approach. In such exploratory settings, our framework remains effective: steps along incorrect reasoning trajectories will exhibit low or negative information gain, indicating they do not contribute meaningfully toward the correct final answer. When the model identifies a more promising path and self-corrects, subsequent steps will show positive information gain, signaling productive progress.
This adaptability to varied reasoning structures is empirically demonstrated in our experiments (Section 5), where we analyse reasoning traces from the MATH/PRM dataset. Steps that human annotators labelled as uninformative or irrelevant consistently show low information gain under our metric, while correct and meaningful steps exhibit high information gain. Thus, our method provides reliable step-wise evaluation regardless of whether the reasoning follows a linear chain or involves branched/exploratory patterns.
We further formalize this remark using our framework in Appendix 8.4.
4 Related Works
Evaluation of CoT reasoning   Several recent works propose methodologies for evaluating CoT reasoning [32, 11, 16, 15, 21, 28, 37, 35]. For example, [16] verifies individual steps in a model’s CoT reasoning by generating multiple LLM responses per prompt and comparing correct responses with incorrect ones.
Similarly, [29, 30] use a fine-tuned LLM to decode multiple reasoning paths from each step and check the correctness of these reasoning paths. However, as we show in our experiments, approaches which simply rely on the correctness of the final output are not sound in general and can lead to false positives. Moreover, these solutions may not be plausible for problems of high difficulty where correct LLM responses might be scarce.
Formalising CoT framework    The formalisation of LLM reasoning remains an active area of research. Most notably [10] introduces a formal framework for LLMs and is a key source of inspiration behind our formalism. Additionally, [9] theoretically examines the expressivity of LLMs with CoT in solving mathematical and decision-making problems, focusing on the transformer architecture’s implications on accuracy. Besides this, [36] provides a formal definition of hallucinations, but does not consider CoT reasoning specifically.
Reward modelling    Outcome-based reward models (ORM) [6, 11, 18] predict the probability of reaching the correct final answer based on a model’s intermediate CoT steps. While they avoid requiring correct intermediate demonstrations, we show in Section 5 that they are unsound for detecting CoT reasoning errors. Step-wise ORM (SORM) [11] extends ORM by estimating the probability of an ‘optimal’ model reaching a correct answer but requires training a larger, more capable model than the base model.
Process-based reward modelling (PRMs) [18, 27] is an alternative approach which directly predicts the correctness of intermediate CoT reasoning steps. Likewise, various other approaches rely on annotated CoT datasets for benchmarking [14, 37, 1, 19, 34, 21, 35, 20]. While these benchmarks and methodologies aid LLM reasoning, collecting annotated data is costly and not easily scalable. In contrast, our approach evaluates an LLM’s CoT reasoning without human-annotated CoT data.
5 Experiments
In this section, we demonstrate our framework’s utility, dubbed Information-Gain (IG) and compare against two baselines for detecting errors in a model’s CoT reasoning. Here we assume access only to the model’s CoT generations $X_{0},X^{M}_{1},\ldots,X^{M}_{T}$ and correct final answers $Y$.
Outcome Reward Model (ORM) [11]
This involves training a classifier, denoted as $f_{\textup{ORM}}$, which takes as input model generations up to any step $t$ in its CoT reasoning, $X^{M}_{t}$, and predicts the probability of the model’s final answer being correct, i.e.
$f_{\textup{ORM}}(X^{M}_{t})\approx\mathbb{P}(Y^{M}=Y\mid X^{M}_{t}).$ (6)
Here, if we observe that this probability of correctness drops significantly after step $t$, i.e. $f_{\textup{ORM}}(X^{M}_{t})\gg f_{\textup{ORM}}(X^{M}_{t+1})$, this indicates that the model applies task $\lambda_{t+1}$ incorrectly.
Math-Shepherd (MS) [29]
This method quantifies the potential for a given reasoning process $X^{M}_{t}$ by using a ‘completer’ model to generate $N$ completions of each reasoning process starting from step $t$, $\{(X^{M}_{t},X^{M}_{t+1,j},\ldots,X^{M}_{T,j},Y^{M}_{j})\}_{j\leq N}$, where $Y^{M}_{j}$ denotes the final answer reached in the $j$’th completion. Then, we estimate the potential of this step based on the proportion of correct answers among the $N$ completions:
$f_{\textup{MS}}(X^{M}_{t})\coloneqq\sum_{j=1}^{N}{\mathbbm{1}}(Y^{M}_{j}=Y)/N.$ (7)
For a fair comparison, we do not assume access to a ‘verifier’ model more capable than our base model. Therefore, we use the base model as the completer model in our experiments.
5.1 Toy data experiments
We first consider a toy setting where we control model behaviour across tasks. Prompts consist of an integer vector $Z_{0}\in\mathbb{Z}^{5}$ sampled from a given distribution. The task $\lambda$ comprises five steps, $\lambda=\lambda_{1}\circ\ldots\circ\lambda_{5}$, where each sub-task $\lambda_{i}$ transforms $Z_{i-1}\in\mathbb{Z}^{5}$ into $Z_{i}\in\mathbb{Z}^{5}$. The correct final answer $Y$ is $Z_{5}$. Additional details on data generation and sub-tasks are in Appendix 9.1.
Generating the dataset    To investigate partial unidentifiability for a given task $\lambda_{i}$ we modify the obtained dataset by introducing ‘noise’ at step $i$. In other words, the task $\lambda_{i}$ is applied incorrectly on a subset of the data, whereas all other tasks are always applied correctly. This represents a model which sometimes fails at step $i$ and we use ‘LLMi’ to denote this model in this experiment. We repeat this procedure for all tasks $\lambda_{i}$ for $i\in\{1,\ldots,5\}$ which yields 5 LLMs $\{\textup{LLM}_{1},\ldots,\textup{LLM}_{5}\}$.
To assess robustness, we introduce a special case in $\textup{LLM}_{3}$, where task $\lambda_{3}$ is applied incorrectly iff the output after $\lambda_{2}$ falls in a set $\mathcal{S}$. This deliberate choice highlights a pitfall of existing baselines and contrasts with other LLMs, where errors occur randomly. In other words, $\lambda_{3}$’s correctness depends on $\lambda_{2}$’s output. For details, see Appendix 9.1.2.
5.1.1 Results
Figure 3 shows how the different baselines quantify the correctness of the different tasks for the 5 different LLMs under consideration. This figure only considers samples where the final answer of the LLM was incorrect, i.e. $Y^{M}\neq Y$. For our method (IG), Figure 3(a) shows the information-gain across the different steps for each LLM. Likewise, Figure 3(b) presents the results for ORM and shows how the average probability of correctness in (6) changes across the different steps, whereas, for Math-Shepherd, Figure 3(c) shows the proportion of correct completions starting after each step (7). Here, any significant drop in the plotted values indicate an incorrect application of a task.
Information-gain rightly quantifies step-wise correctness     We observe that the information-gain remains positive for each LLM until we encounter an incorrect reasoning step, at which point it drops to negative values. Therefore, our method can identify the incorrectly executed task for each LLM under consideration. We used a GPT-2 supervisor model to estimate information-gain.
Pitfall of the baselines     While ORM and Math-Shepherd usually identify incorrect reasoning steps, they fail for $\textup{LLM}_{3}$. This is because $\lambda_{3}$ is misapplied iff the output after $\lambda_{2}$ lies in $\mathcal{S}$. Thus, the classifier can predict the final output’s correctness at $\lambda_{2}$ by checking if $Z_{2}$ lies in $\mathcal{S}$, leading to overconfidence in error detection at $\lambda_{2}$ instead of $\lambda_{3}$.
(a) Our results
(b) ORM results
(c) Math-Shepherd results
Figure 3: Heatmaps (left) quantifying the correctness of different sub-tasks for the 5 LLMs using the different baselines, and the associated classification metrics (right). Red color in the heatmaps indicates a significant drop in the plotted metrics (an incorrectly executed sub-task).
Similarly, with Math-Shepherd for $\textup{LLM}_{3}$ (using the same model as a completer), a completion is incorrect if the output after $\lambda_{2}$ lies in $\mathcal{S}$. Here, all completions fail, regardless of the starting step, making it impossible to pinpoint where $\textup{LLM}_{3}$ goes wrong.
Table 1: Sample-wise classification of a sub-task for $\textup{LLM}_{3}$.
Acc $\uparrow$ TPR $\uparrow$ FPR $\downarrow$
IG (Ours) 0.96 0.98 0.06
ORM 0.77 0.98 0.54
MS 0.60 1.0 1.0
Sample-wise detection     We also use the different baselines for sample-wise detection of erroneous steps, as outlined in Section 3.3. A step is classified as incorrect if a baseline’s metric falls below a threshold. Table 1 presents the results for $\textup{LLM}_{3}$, with optimal thresholds chosen from a held-out dataset. Our method achieves significantly higher accuracy and fewer false positives than the baselines, making it more reliable for sample-wise error detection.
5.2 Arithmetic operations on Llama-3-8B
Following our toy experiments, we now evaluate our framework in a more realistic setting using the Llama-3-8B model [7]. We focus on a simple arithmetic task that involves both multiplication and addition. The goal is to assess the model’s performance on each operation.
Experimental setup     We sample two integers $x$ and $y$ uniformly from $[1,10^{5})$. The prompt to the model is:
Prompt:x = {x}, y = {y}, Please calculate the following: 1. 3x,    2. 2y,    3. 3x + 2y
Model Accuracy     The model’s accuracy across steps is:
Step 1: 80%, Step 2: 98%, Step 3: 42%.
Most failures occur in Step 3, which involves adding previously computed values. Analyzing the $(x,y)$ distribution where the model is incorrect (Figure 4), we find that errors mainly arise when one variable is large and the other is small. This suggests that correctness depends heavily on $(x,y)$, making it difficult for baselines to pinpoint the erroneous step in the model’s CoT reasoning.
Figure 4: The distribution of $(x,y)$ for incorrect samples: Llama-3-8B struggles to add large and small numbers (represented by the top-left and bottom-right shaded regions).
5.2.1 Results
Our method     We trained the supervisor model by fine-tuning a Llama-3-8b model using Low Rank Adaptation (LoRA) [13]. Table 2 shows that there is a significant drop in information-gain at step 3 relative to steps 1 and 2, demonstrating that our method correctly identifies that the failure mainly occurs at step 3.
Outcome Reward Model (ORM)     For ORM, the mean probability of correctness (Table 2) remains unchanged at each step. Figure 4 suggests this is because ORM predicts correctness based solely on $x$ and $y$ in the prompt. Crucially, its confidence remains constant even as intermediate reasoning steps are added, preventing it from distinguishing the model’s performance at different steps.
Math-Shepherd (MS)     Table 2 shows the proportion of correct completions for MS. While this is low at step 3, only 5-7% of completions from steps 1 and 2 yield a correct output, despite errors mostly occurring at step 3. This is because Llama-3-8B’s correctness is largely determined by $(x,y)$ in the prompt (Figure 4). As a result, MS frequently mislabels steps 1 and 2 as incorrect, leading to a higher false positive rate compared to our baseline.
Table 2: Experimental results for Toy Arithmetic, GSM8K and PRM800K experiments. In each of the datasets, we denote the “correct" and “wrong" steps with and respectively.
Datasets & Methods Operations Sample-wise detection metrics
Toy Arithmetic Methods Step 1: $3x$ Step 2: $2y$ Step 3: $3x+2y$ Accuracy $\uparrow$ TPR $\uparrow$ FPR $\downarrow$
IG (Ours) 0.67 0.24 0.027 - 0.76 0.51 0.02
ORM 0.24 0.24 0.24 - 0.56 0.10 0.07
MS 0.068 0.059 0.00069 - 0.53 0.99 0.86
GSM8K Methods Addition Multiplication Division Subtraction Accuracy $\uparrow$ TPR $\uparrow$ FPR $\downarrow$
IG (Ours) 0.99 0.026 1.05 1.06 0.72 0.95 0.62
ORM 0.46 0.024 0.38 0.013 0.58 1.00 1.00
PRM800K Methods Negative Neutral Positive - Accuracy $\uparrow$ TPR $\uparrow$ FPR $\downarrow$
IG (Ours) 0.058 -0.011 0.168 - 0.74 0.84 0.37
ORM 0.734 0.745 0.744 - 0.69 0.55 0.18
Sample-wise detection     When using these methods for sample-wise detection of incorrect steps, our approach yields the highest accuracy among the baselines considered. This superior performance is attributed to the fact that baselines like ORM and MS often falsely flag steps 1 and 2 as incorrect, as evidenced by their high FPRs in Table 2.
5.3 Experiments on the controlled GSM8K Dataset
To evaluate our method on a complex dataset, we conducted experiments on GSM8K [6], controlling specific factors for more interpretable results.
We begin by using GPT-4 [23] to generate answers for GSM8K questions where the “multiplication” operation is always done incorrectly, while all other operations are correct. Next, we filtered the dataset to ensure that “multiplication”, “subtraction”, and “addition” never appeared together within the same Chain of Thought (CoT) solution. In particular, we ensured in our setting that, all incorrect final answers included both “multiplication” and “subtraction”, whereas correct final answers did not involve either operation. This introduces a spurious correlation between “subtraction” and wrong answers.
In this setup, we mainly focused on evaluating ORM and our proposed method, as MS (with the same completer) fails trivially under these conditions. Specifically, “multiplication” is inherently unidentifiable, since any CoT containing “multiplication” negates the influence of other sub-tasks by design. Further details on the experimental setup can be found in Appendix 9.3.
5.3.1 Results
Table 2 shows that our information-theoretic approach (IG) successfully identifies the unidentifiable sub-task. Since the “multiplication” rule is intentionally incorrect, it yields minimal to no information gain, as expected. However, ORM results reveal a different pattern: both “multiplication” and “subtraction” have low correctness probabilities, as they are linked to incorrect final answers. This suggests that the standard ORM approach may misleadingly classify “subtraction” as incorrect.
Additionally, in our sample-wise experiment, we observe a similar trend when we use the methods to assess the sample-wise correctness of “multiplication" and “subtraction" for each prompt. Here, our proposed method not only accurately detects the unidentifiable sub-task but also highlights a significant shortcoming of ORM. Specifically, ORM falsely flags “subtraction”, which is actually correct, as an incorrect sub-task due to spurious correlations.
5.4 Experiments on the PRM800K dataset
To further demonstrate the practical applicability of our method, we have conducted an additional experiment on OpenAI’s PRM800k dataset [18] which is obtained by labeling the intermediate steps of the MATH dataset [12].
More specifically, this dataset is a process supervision dataset with step-level correctness labels for model-generated solutions to MATH problems. To create it, Lightman et al. [18] asked human annotators to label each step from fine-tuned GPT-4 solutions as positive (+1), negative (-1), or neutral (0). A positive label indicates a correct, reasonable step; a negative label denotes an incorrect or unreasonable step; and a neutral label indicates ambiguity (e.g., subtly misleading or technically valid yet poor).
The objective is to identify incorrect Chain-of-Thought (CoT) steps, specifically those labelled as (-1) by annotators, using our method as well as ORM. However, we do not utilize the step-wise labels during the process; they are only used for evaluation purposes. Since the base GPT-4 model used to generate the PRM data is not publicly available, we were unable to obtain MS results for this dataset.
5.4.1 Results
Table 2 shows the information gain and mean correctness probability for positive, negative, and neutral sub-steps.
As expected, these results show that the information-gain is significantly lower for incorrect steps (with labels -1) than for labels +1. Additionally, we also observe that the information-gain is negative for neutral steps (with labels 0), which is explained by the fact that these steps do not add any useful information regarding the ground truth (as these were deemed irrelevant/ambiguous by the human labellers). In contrast, the average probability of correctness for the ORM classifier is roughly the same across each label and, on average, is not very informative.
Sample-wise detection    Additionally, we also used the sample-wise information-gain (IG) as well as the ORM baseline to classify if a step is correct (as outlined in Section 3.3). To avoid ambiguity, we filtered out the neutral sub-steps (with labels 0) for this experiment and considered a balanced held-out dataset with equal number of correct and incorrect steps. Table 2 also shows the sample-wise results for both methods (where we chose the best thresholds for each baseline using a held-out dataset).
It can be seen that the accuracy of our method is higher than that of the ORM classifier. Additionally, our method also leads to higher TPR (and hence a lower FNR) than the ORM classifier. These results show that our method outperforms the outcome-based baselines on more complex datasets such as the MATH data as well.
6 Discussion and Limitations
In this paper, we introduce a novel information-theoretic approach to evaluate Chain-of-Thought (CoT) reasoning in LLMs without annotated intermediate steps. Our framework effectively identifies erroneous reasoning across diverse settings and consistently outperforms baselines, including Outcome Reward Models (ORMs) [11] and Math-Shepherd (MS) [29]. However, our approach does have some limitations.
Although our method avoids human-annotated step-wise data, it requires additional training of the supervisor model, which is computationally expensive. Future work could explore in-context learning to estimate information gain, reducing training needs and improving efficiency. Additionally, while our method does not require correctness labels for every step, we still need to categorize each step according to its respective sub-task. However, this limitation is not unique to our method, as both ORM and MS also rely on such labels to draw sub-task-specific conclusions.
Lastly, while we focus on logical and mathematical datasets, our method also extends to other domains requiring CoT reasoning, such as Blocks World [25]. As we discuss in Appendix 8.2, this is an interesting avenue which we leave for future research.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Acknowledgements
We would like to thank Li Hang for his valuable insights and guidance during the development of this work.